OpenStudy (lichking):

One day, a person went to a horse racing area. Instead of counting the number of humans an horses, he counted 74 heads and 196 legs. How many humans and horses were there?

OpenStudy (campbell_st):

so let y = humans and x = horses so for heads x + y = 74 and legs 4x + 2y = 196 which could be written 2x + y = 98. now all you need to do is solve simultaneously.... either by elimination or substitution. hope he helps

OpenStudy (lichking):

Aye it is helpful,so it is